OpenStudy (anonymous):

For a motorcyle traveling at speed v(in mph) when the brakes are applied the distance d in (feet) required to stop the motorcycle may be approximated by the formula d=0.05v^2. Find the instantenous rate of change of distance with the respect of velocity when the speed is 40 mph

OpenStudy (anonymous):

instantenous rate of change of w.r.t of velocity = d ( d(v))/dv = 2*v/20 = v/10 = 40/10 = 4 unit

OpenStudy (anonymous):

ok got it thank u:)